OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9)

OpenStudy (anonymous):

slope m=y2-y1.x2-x1=9-6/2-1=3 y-6=3(x-1) y-6=3x-3 y=3x+3

OpenStudy (cwrw238):

slope = (9-6)/2-1) = 3 use the general form y-y1 = m(x-x1) where m = slope = 3, and (x1, y1) is say (1,6) so y - 6 = 3(x - 1) y = 3x - 3 + 6 y = 3x + 3

OpenStudy (cwrw238):

lol - i don't why i bother when sai1234 is around - he's like lightning

OpenStudy (cwrw238):

do you follow that ok john?

OpenStudy (anonymous):

(1,6) (x1, y1) (2,9) ( x2, y2) M(slope) = y2- y1/ x2-x1                   = 9-6/ 2-1                 =3 Equation in intercept form is y- y1 = m ( x-x1) Substituting we get y- 6 = 3 ( x-1) y= 3x+3